OpenStudy (anonymous):

NEED HELP WITH ALGEBRA 1 PLEASE HELP :( :(

OpenStudy (anonymous):

What do you need help with?

OpenStudy (anonymous):

A box has dimensions of 19 inches long, 1.7 feet wide, and 6 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. A. 193.8 cubic inches (in3) B. 2325.6 cubic inches (in3) C. 1.35 cubic inches (in3) D. 16.15 cubic inches (in3)

OpenStudy (anonymous):

I've messaged you explaining it but the answer is 'A', 193.8in^3

OpenStudy (texaschic101):

just multiply them V = L * W * H V = 19 * 1.7 * 6 V = 193.8 in^3

OpenStudy (anonymous):

Oh ok thank u